OpenStudy (anonymous):

Write The Trig Expression As An Algebraic Expression cos(2arcsin(x)) I can write it fine with normal angles, but I don't know if I should let the arbitrary variable be equal to arcsin(x), 2arcsin(x), or something over 2. The answer should be 1-2x^2

OpenStudy (tkhunny):

\(\cos(2\alpha) = 1 - 2\sin^{2}(\alpha)\)